# Bad Students Make Great Teachers: Active Learning Accelerates Large-Scale Visual Understanding

### v2

@@ A.1 Details of total compute calculations
$C_{\text{easy\_ref}}=\big{(}3F_{\text{learn}}+\rho F_{\text{ref}}\big{)}\beta+3F_{\text{ref}}$
where $\rho=B/b$ (set to $\rho=2.0$ for our experiments) and $\beta$ is the learner speedup vs. IID. The average FLOPs per learner update for RHO is:
$C_{\text{RHO}}=\big{(}3F_{\text{learn}}+\rho(F_{\text{learn}}+F_{\text{ref}})\big{)}\beta+3F_{\text{ref}}$
And for ActiveCLIP / ClassAct:
$C_{\text{ClassAct / ActiveCLIP}}=\big{(}3F_{\text{learn}}+2\rho F_{\text{ref}}\big{)}\beta+3F_{\text{ref}}$
@@
$\Big{(}3F_{\text{learn}}+\rho F_{\text{act}}\Big{)}\beta+3F_{\text{ref}}<3F_{\text{learn}}$
where $F_{\text{act}}$ is one of $F_{\text{ref}}$, $\big{(}F_{\text{learn}}+F_{\text{ref}}\big{)}$ or $2F_{\text{ref}}$ for “easy reference”, RHO or ClassAct / ActiveCLIP, respectively.
A.2 Implementation details: ClassAct
@@
Architectures: We use standard vision transformers trained with 224$\times$224 image resolution and a patch size of 16$\times$16. We consider the standard ViT-Base and ViT-Large variants as learners in our in our main study, and smaller-scale variants (ViT-S and ViT-Tiny) when reducing the scoring-model computation. We introduce even smaller variants to scale down actor computation further, and detail the configurations of this “ViT-Micro” family in section A.7.
@@
Evaluation: We evaluate the performance of the learned model by applying it to a held-out set of JFT images, and measuring top-1 accuracy. Previous work has found that in-domain performance on JFT is highly predictive of out-of-distribution generalization [44], hence we leave the study of transfer performance to the multimodal setting (Sec. A.3).
A.3 Implementation details: ActiveCLIP
@@
A.4 Online ClassAct / ClassCLIP
ClassAct / ClassCLIP can also be run in a single training loop, by training the reference model on the super-batch:
@@ A.4 Online ClassAct / ClassCLIP
3:    $X\sim\mathcal{D}$, where $|X|=B$triangleright Sample IID
@@
We found that $B=10b$ was sufficient to reproduce the pre-trained reference (Algorithm 2) case where $B=2b$, as described Section 4.4. It is possible that the super-batch size could be shrunk, but we did not evaluate our method in between these values.
A.5 Analyzing learnability scores

### v4

@@ 0.A.1 Details of total compute calculations
$C_{\text{easy\_ref}}=\big{(}3F_{\text{learn}}+\frac{F_{\text{ref}}}{\text{SPI}}\big{)}\beta+3F_{\text{ref}}$
where $\text{SPI}=b/B$ ($=0.5$ for our experiments) and $\beta$ is the learner speedup vs. uniform sampling. The average FLOPs per learner update for RHO is:
$C_{\text{RHO}}=\big{(}3F_{\text{learn}}+(F_{\text{learn}}+\frac{F_{\text{ref}})}{\text{SPI}}+\big{)}\beta+3F_{\text{ref}}$
And the average FLOPs per update for ActiveCLIP / ClassAct is:
$C_{\text{ClassAct / ActiveCLIP}}=\big{(}3F_{\text{learn}}+2\frac{F_{\text{ref}}}{\text{SPI}}\big{)}\beta+3F_{\text{ref}}$
@@
$\Big{(}3F_{\text{learn}}+\frac{F_{\text{act}}}{\text{SPI}}\Big{)}\beta+3F_{\text{ref}}<3F_{\text{learn}}$
where $F_{\text{act}}$ is one of $F_{\text{ref}}$, $\big{(}F_{\text{learn}}+F_{\text{ref}}\big{)}$ or $2F_{\text{ref}}$ for "easy reference", RHO or ClassAct / ActiveCLIP, respectively.
0.A.2 Implementation details: ClassAct
@@
Architectures: We use standard vision transformers trained with 224$\times$224 image resolution and a patch size of 16$\times$16. We consider the standard ViT-Base and ViT-Large variants as learners in our in our main study, and smaller-scale variants (ViT-S and ViT-Tiny) when reducing the scoring-model computation. We introduce even smaller variants to scale down actor computation further, and detail the configurations of this “ViT-Micro” family in section 0.A.7.
@@
Evaluation: We evaluate the performance of the learned model by applying it to a held-out set of JFT images, and measuring top-1 accuracy. Previous work has found that in-domain performance on JFT is highly predictive of out-of-distribution generalization [43], hence we leave the study of transfer performance to the multimodal setting (Sec. 0.A.3).
0.A.3 Implementation details: ActiveCLIP
@@
0.A.4 Online ClassAct / ActiveCLIP
ClassAct / ActiveCLIP can also be run in a single training loop, by training the reference model on the super-batch:
@@ 0.A.4 Online ClassAct / ActiveCLIP
3:    $X\sim\mathcal{D}$, where $|X|=B$triangleright Sample uniformly
@@
We found that $B=10b$ (SPI=0.1) was sufficient to reproduce the pre-trained reference (Algorithm 1) case where $B=2b$ (SPI=0.5), as described Section 4.4. It is possible that the super-batch size could be shrunk, but we did not evaluate SPI values in between these values.
0.A.5 Analyzing learnability scores
@@ 0.A.6 Learning infrastructure
This setup allows us to run ActiveCLIP at the same speed as uniform sampling experiments, even without approximating the scoring models. This distributed setup may be advantageous over the sequential online batch selection setup (where the actors and learners run on the same hardware) in cases where time, not compute, is the bottleneck; adding computation to parallelize the data filtration is an alternative to increasing the batch size, which is known to saturate [44].
@@ 0.A.7 Small ViT models
Appendix 0.B Appendix: supplementary results
0.B.1 FLOP efficient contrastive training
Our previous results (Table 2) showed that it is possible to improve the performance of contrastve pre-training using active data selection via ActiveCLIP. However, these results used a ViT-B policy (reference and online models) to train a ViT-B learner. To explore whether similar gains could be achieved with fewer FLOPs, we repeated out analyses using ViT-Ti policies (Table 5). As in our ClassAct JFT experiments, our results show that performance degrades marginally, eve though the ViT-Ti policy requires 50x fewer FLOPs.
IN-1K COCO
Method Train ex. ZS Top-1 im2txt txt2im
CLIP [30] 13B 68.3 52.4 33.1
EVA-CLIP [37] 3B 69.7
SigLIP [44] 3B 68.8 56.7 40.2
ActiveCLIP-B 3B 70.3 57.8 42.6
ActiveCLIP-Ti 3B 69.9 57.2 42.4
Table 5: Comparison of ActiveCLIP to public multimodal pre-taining methods Extension of Table 2, replacing the ViT-B reference and online models with a ViT-Ti. The ViT-Ti policy marginally underperforms the ViT-B policy, but is also SOTA vs. previous baselines.
We also repeated this ablation for our out-of-domain contrastive pre-training experiments shown in Fig. 6). Again, ViT-Ti policies produces similar speedups and were also compute-positive, mirroring our ClassAct JFT experiments (Figure 5), 4).
ImageNet 0-Shot COCO (im2txt) COCO (txt2im)
Method Ref. Speed FLOP Top 1 Speed FLOP R@1 Speed FLOP R@1
model -up saving Acc. -up saving -up saving
IID 0.0% 0.0% 0.386 0.0% 0.0% 0.256 0.0% 0.0% 0.457
ActiveCLIP ViT-Ti 43.3% 26.2% 0.428 38.7% 20.8% 0.292 38.3% 20.3% 0.509
ActiveCLIP ViT-B 51.0% -163.4% 0.449 48.3% -172.4% 0.305 47.3% -175.6% 0.532
Table 6: Compute-positive contrastive pre-training. Reproduction of experiments from Figure 6, replacing the ViT-B reference and online models with ViT-Ti. As in our ClassAct JFT experiments (Figure 5), 4), the ViT-Ti policy produced nearly the same learner speedups as the ViT-B policy, but unlike the ViT-B policy also produced net FLOP savings. In all cases the reference model is trained on LTIP and the learner is trained on ALIGN.